Urn One has 11 red balls and 8 blue balls. Urn Two has 10 red balls and 5 blue balls. An urn is selected at random and then a ball is selected from the selected urn at random. What is the probability that the ball selected is blue? State your answer to three places of decimal.

Your Answer:
Answer #1

P(Blue in urn 1) = 8 / (11+8) = 8/19

P(Blue in urn 2) = 5 / (10+5) = 5/15

An urn is selected at random, so probability of selection of urn = 1/2

P(Blue balls) = 1/2 [ P(Blue in urn 1) + P(Blue in urn 2) ]

= 1/2 [ 8/19 + 5/15 ]

= 0.377